Solve the following equation by using the quadratic formula. (Enter solutions from smallest to largest. If there is one solution with a multiplicity of two, enter it in both answer boxes.)
x^2-16x+64=0
x = ?
x = ?
Please show me the steps to get the answer. Thanks

Answer by stanbon(75887)

Solve the following equation by using the quadratic formula. (Enter solutions from smallest to largest. If there is one solution with a multiplicity of two, enter it in both answer boxes.)
x^2-16x+64=0
Factor:
x^2-8x-8x+64 = 0
x(x-8)-8(x-8) = 0
(x-8)(x-8) = 0
--------------------
x = 8 with multiplicity 2.
==============================
Cheers,
Stan H.
====================